3x - y = 6
3x - y)/3 = 2
A(4, -1)
B(1, -2)
C(-5, 2)
D(-1, 5)
ENo Solution
F3x - 6
I'm really confused on this one, I'm not sure what to do. If someone could show me the steps and the correct answer I would appericiate it very much. Thank you for your time!

Answer by Alan3354(69443)

3x - y = 6
3x - y)/3 = 2
A(4, -1)
B(1, -2)
C(-5, 2)
D(-1, 5)
ENo Solution
F3x - 6
I'm really confused on this one, I'm not sure what to do. If someone could show me the steps and the correct answer I would appericiate it very much. Thank you for your time!
---------------
3x - y = 6
3x - y)/3 = 2
-----------
I think for the 2nd eqn you meant: (3x - y)/3 = 2, there's no left parenthesis.
If so, then the 2 eqns are the same, the 2nd is the 1st divided by 3. So it's one equation in 2 unknowns which means there's no solution.
E No solution